### v1

# Flag complex face structures and decompositions

Abstract
We consider how $f$-vectors of Cohen–Macaulay and vertex decomposable flag complexes in geometric settings decompose and interact with geometric transformations. This includes subdivisions of simplicial complexes and explicit equations connecting them to $f$-vectors. Applications from this approach include positivity properties of polynomials built out of simplicial complexes lying between unimodality and real-rootedness.
Introduction
While the $h$-vectors of simplicial complexes have many well-known algebraic and geometric structures, their $f$-vectors often do not readily admit structural properties that apply in general. We use the case of flag simplicial complexes satisfying natural algebraic properties to study how they decompose and interact with geometric transformations. Along the way, we will see connections to positivity questions on reciprocal/palindromic polynomials.
The first structural property we will consider is Boolean decompositions of simplicial complexes (Definition 1.2), which come up in the context of the positivity questions mentioned above. The specific motivating result is the following:
Proposition 0.1.
(Nevo–Petersen–Tenner, Proposition 6.2 on p. 1377 of [24])
Let $\Delta$ be a flag sphere. If $\gamma(\Delta)$ is the $f$-vector of a simplicial complex $S$, then $h(\Delta)$ is the $f$-vector of the following simplicial complex:
$\Gamma=\{F\cup G:F\in\mathcal{F}(\gamma(\Delta)),G\in 2^{[d-2|F|]}\},$
where $\mathcal{F}(T)$ is the standard compressed complex associated to $T$ taking $\frac{d}{2}$-compressed complexes $\mathcal{F}_{k}(f(T))$ and letting $k\to\infty$ (see p. 263 of [34] and p. 1366 – 1369 of [24]).
Note that the part involving $|F|$ is multiplied by $2$.
The positivity question mentioned above is gamma positivity surveyed in [3], which connects a wide variety of topics such as permutation statistics, triangulations of simplicial complexes, and the Hopf conjecture for certain piecewise Euclidean manifolds.
In this direction, we obtain some results on how these decompositions interact with transformations and recursive properties of flag simplicial complexes.
Theorem 0.2.
  1. 1.
    (Proposition 1.5) The edge subdivision of a simplicial complex that has a Boolean decomposition (Definition 1.2) has a generalized Boolean decomposition (Definition 1.4).
  2. 2.
    (Part 1 of Corollary 1.7) Suppose that a Cohen–Macaulay simplicial complex $\Delta$ has an associated simplicial complex $\Gamma$ such that $h(\Delta)=f(\Gamma)$ and $\Gamma$ has a Boolean decomposition. Then, the same is true for an edge subdivision $\Delta^{\prime}$ of $\Delta$. In other words, there is a simplicial complex $\Gamma^{\prime}$ such that $h(\Delta^{\prime})=f(\Gamma^{\prime})$ and $\Gamma^{\prime}$ has a Boolean decomposition.
  3. 3.
    (Part 2 of Corollary 1.7) Suppose that $\Delta$ is a simplicial complex satisfying one of the following properties:
    • $\Delta$ is vertex decomposable and flag.
    • $\Delta$ is a balanced, vertex decomposable, flag simplicial complex.
    Any simplicial complexes associated to antistars and links of vertices with Boolean decompositions can be used to construct a flag (resp. quasi-flag from Definition 4.2 on p. 99 of [10]) simplicial complex $\Gamma$ such that $h(\Delta)=f(\Gamma)$ and $\Gamma$ has a Boolean decomposition.
Next, we consider connections between structures of $f$-vectors and positivity properties of reciprocal/palindromic polynomials. The main tool we use is an interpretation of the gamma polynomial in terms of an inverted Chebyshev expansion which we found in [26]. This was motivated by thinking about gamma polynomial-like invariants of the “usual” basis of reciprocal/palindromic polynomials of a given degree.
Theorem 0.3.
(Gamma vectors and inverted Chebyshev expansions, Theorem 1.7 on p. 7 of [26])
Let $h(t)=h_{0}+h_{1}t+\ldots+h_{d-1}t^{d-1}+h_{d}t^{d}$ be a reciprocal polynomial satisfying the relations $h_{k}=h_{d-k}$ for $0\leq k\leq\frac{d}{2}$. Then, the gamma-polynomial associated to $h(t)$ is equal to
$\gamma(u)=u^{\frac{d}{2}}g\left(\frac{1}{u}-2\right),$
where
$g(u)\coloneq h_{\frac{d}{2}}+2\sum_{j=1}^{\frac{d}{2}}h_{\frac{d}{2}-j}T_{j}\left(\frac{u}{2}\right).$
This is a sort of “inverted Chebyshev basis expansion”.
Alternatively, we can rewrite this as
$u^{\frac{d}{2}}\gamma\left(\frac{1}{u}\right)=g(u-2)$
or
$(u+2)^{\frac{d}{2}}\gamma\left(\frac{1}{u+2}\right)=g(u).$
This enables us to apply the following result of Hetyei on compatibility between a variant of the usual $f$-polynomial and Tchebyshev transforms $T:\mathbb{R}[x]\longrightarrow\mathbb{R}[x]$ sending $x^{n}\mapsto T_{n}(x)$.
Theorem 0.4.
(Hetyei, Proposition 3.3 on p. 578 – 579 of [18])
Given a simplicial complex $A$, let the $F$-polynomial of $A$ be $F(x)\coloneq f\left(\frac{x-1}{2}\right)$. The Tchebyshev transform of the $F$-polynomial of $A$ is the $F$-polynomial of its Tchebyshev triangulation (ordered repeated edge subdivisions):
$T(F_{A})(x)=F_{T(A)}(x).$
As a result, we find a connection with signed unused color complexes (Definition 2.19) of balanced simplicial complexes (Theorem 2.22).
Theorem 0.5.
Let $\Delta$ be a flag simplicial homology sphere of even dimension $d$ and
$P_{\Delta}(u)\coloneq h_{\frac{d}{2}}(\Delta)+2\sum_{j=1}^{\frac{d}{2}}h_{\frac{d}{2}-j}(\Delta)u^{j},$
which is $g(2u)$ with $T_{k}(u)$ replaced by $u^{k}$ (i.e. $g(2u)=T(P(u))$ and $P(u)=T^{-1}g(2u)$).
Given $k\geq 1$, there is a $\left(\frac{d}{2}-1\right)$-dimensional balanced simplicial complex $\Gamma$ such that
$[\alpha^{k}]\frac{P_{\Delta}(2\alpha+1)+f_{\frac{d}{2}-1}(\Gamma)}{2}-\sum_{j=0}^{\frac{d}{2}}f_{\frac{d}{2}-j-1}(\Gamma)+1=f_{k-1}(\mathcal{D}(\Gamma))$
and
$\frac{P_{\Delta}(2\alpha+1)+f_{\frac{d}{2}-1}(\Gamma)}{2}-\sum_{j=0}^{\frac{d}{2}}f_{\frac{d}{2}-j-1}(\Gamma)+1=f_{\mathcal{D}(\Gamma)}(\alpha).$
Substituting in $\alpha=\frac{\beta-1}{2}$, we have that
$\frac{P_{\Delta}(\beta)+f_{\frac{d}{2}-1}(\Gamma)}{2}-\sum_{j=0}^{\frac{d}{2}}f_{\frac{d}{2}-j-1}(\Gamma)+1=F_{\mathcal{D}(\Gamma)}(\beta),$
where
$F(u)\coloneq\sum_{i=0}^{D}f_{i-1}(A)\left(\frac{u-1}{2}\right)^{i}$
for a cell complex $A$ of dimension $D-1$.
Combining this with Theorem 0.4, we find that the the $f$-polynomial of the interval poset of $S$ is a scaling of that of the mirroring/Danzer complex/power complex of the Tchebyshev subdivision of some simplicial complex associated to $\Gamma$ (Corollary 2.33).
We start by directly studying connection between gamma vectors and flag spheres in Section 1. This involves an overview of Boolean decompositions and gamma positivity followed by an analysis of its behavior under edge subdivisions (Proposition 1.5) and recursive properties connected to vertex decomposability (Corollary 1.7). Afterwards, we set up background information for analyzing related geometry of the simplicial complexes giving the relevant $f$-vectors in Section 2.1. A transformation that will occur frequently is taking the poset of intervals partially ordered by inclusion in Section 2.2. Putting this together, we find a connection between Tchebyshev subdivisions (repeated edge subdivisions) of “unused coloring complexes” (Definition 2.19) and interval posets of $S$ such that $\gamma(\Delta)=f(S)$ (Corollary 2.33). Finally, we use generalizations of Tchebyshev subdivisions to cell complexes and the idea behind a proof of a result of Rowlands to study connections between $f$-vectors of flag complexes coming from gamma vector constructions and $CAT(0)$ complexes.
1 Boolean decompositions of face vectors of simplicial complexes
A property that was studied in the context of possible underlying structures for gamma positivity is symmetric Boolean decompositions (see Section 4.7 on p. 84 – 85 of [27]). Another instance where Boolean decompositions appear is in possible structures of simplicial complexes whose $f$-vectors are $h$-vectors of gamma-nonnegative flag spheres.
Proposition 1.1.
(Nevo–Petersen–Tenner, Proposition 6.2 on p. 1377 of [24])
If $\gamma(\Delta)$ is the $f$-vector of a simplicial complex $S$, then $h(\Delta)$ is the $f$-vector of the following simplicial complex:
$\Gamma=\{F\cup G:F\in\mathcal{F}(\gamma(\Delta)),G\in 2^{[d-2|F|]}\},$
where $\mathcal{F}(T)$ is the standard compressed complex associated to $T$ taking $\frac{d}{2}$-compressed complexes $\mathcal{F}_{k}(f(T))$ and letting $k\to\infty$ (see p. 263 of [34] and p. 1366 – 1369 of [24]).
Note that the part involving $|F|$ is multiplied by $2$.
Since the identity expressing $h_{i}$ as a linear combination of $\gamma_{j}$ used to prove the result above (Observation 6.1 on p. 1377 of [24]) is an invertible linear transformation, having such a decomposition for the simplicial complex $\Gamma$ such that $h(\Delta)=f(\Gamma)$ (which exists by work of Caviglia–Constantinescu–Varbaro [8] and Björner–Frankl–Stanley [6]) is equivalent to the gamma vector of $\Delta$ being the $f$-vector of some simplicial complex. This structure can be recorded in a definition.
Definition 1.2.
A simplicial complex $\Gamma$ of dimension $d-1$ has a Boolean decomposition if it an be expressed as
$\Gamma=\{F\cup G:F\in S,G\in 2^{[d-2|F|]}\}$
for some simplicial complex $S$.
We study how this decomposition behaves under transformations often used on flag simplicial complexes applied to $\Gamma$ or $\Delta$. The result below is one indication that the latter is also natural to consider.
Theorem 1.3.
(Constantinescu–Varbaro, Theorem 4.3 on p. 99 of [10])
Let $\Delta$ be a balanced, vertex decomposable, flag simplicial complex on $[n]$. Then there exists a quasi-flag simplicial complex $\Gamma$ such that $f(\Gamma)=h(\Delta)$.
In addition, the $f$-vector of a flag simplicial complex $\Gamma$ is the $h$-vector of some vertex decomposable flag complex $\Delta$ by work of Cook–Nagel (Corollary 3.10 on p. 94 of [11]). Conjecturally, the $h$-vector of every vertex decomposable flag complex is the $f$-vector of some flag complex (Corollary 3.11 on p. 94 of [11]). We consider how possible structures on $\Gamma$ compare with those of other flag simplicial complexes that are PL homeomorphic to it. By a result of Lutz–Nevo (Theorem 1.2 on p. 70 and p. 77 of [21]), it suffices to consider edge subdivisions and contractions of $\Gamma$.
This motivates a generalization of the Boolean decompositions.
Definition 1.4.
A generalized Boolean decomposition of a simplicial complex $\Gamma$ of dimension $d-1$ is a decomposition of $\Gamma$ as the disjoint union of collections of faces of the form
$\{F\cup G:F\in S,G\subset 2^{[d-2|F|]}\}$
where $S$ is either a simplicial complex or the open star of one and $G$ is a subcollection of the $2^{[d-2|F|]}$ restricted by size or containing a specific element (similar to the open star).
We use this to show the following:
Proposition 1.5.
The edge subdivision of a simplicial complex that has a Boolean decomposition has a generalized Boolean decomposition.
Proof.
Let $d=\dim\Gamma+1$ be an even number. Suppose there is some simplicial complex $S$ of dimension $\frac{d}{2}-1$ such that $\Gamma=\{F\cup G:F\in S,G\in 2^{[d-2|F|]}\}$. Let $e=\{a,b\}\in\Gamma$ be an edge of $\Gamma$. We consider where there is such a decomposition for the edge subdivision $\Gamma^{\prime}$ of $\Gamma$ with respect to $e$. Let $v\in V(\Gamma^{\prime})$ be the vertex subdividing what was the edge $e$ in $\Gamma$. In general, we have that $\operatorname{ast}_{\Gamma^{\prime}}(v)\cong\operatorname{ast}_{\Gamma}(e)$ while faces of $\Gamma$ that contained $e$ are doubled by $v$ and $\operatorname{lk}_{\Gamma}(e)\cong\operatorname{lk}_{\Gamma^{\prime}}(a,v)\cong\operatorname{lk}_{\Gamma^{\prime}}(b,v)$ (with $v$ replacing one of $a$ or $b$). We split this into 3 cases:
  1. 1.
    Suppose that $a,b\in[d]$ (i.e. the Boolean part $G$ of the decomposition). Without loss of generality, we will assume that $a<b$. Note that there is an extra new vertex $v$ involved in the “Boolean” part. Then, $S$ and faces of the form $F\cup G$ with $d-2|F|\leq a-1$ or $d-2|F|\geq b+1$ (equivalently $|F|\geq\frac{d-a+1}{2}$ or $|F|\leq\frac{d-b-1}{2}$) in general are unaffected. This means that the faces $F\in S$ to consider are those with
    $\frac{d-b}{2}\leq|F|\leq\frac{d-a}{2}.$
    These are faces which involve neither $a$ nor $b$. In general, those containing only one of $a$ or $b$ are also unaffected. The ones that used to contain both $a$ and $b$ are “doubled” with this pair replaced by $a$ and $v$ or $b$ and $v$. We continue to use the same faces $F\in S$ while subsets of $[d-2|F|]$ (equivalently of $[d]$ in general) containing both $a$ and $b$ are doubled. These come from $F\in S$ such that $b\leq d-2|F|$ (equivalently $|F|\geq\frac{d-b}{2}$). While there is a decomposition into two parts, the decomposition is different being a purely Boolean decomposition. That being said, the added part has a decomposition into two parts with the first part from $F\in S$ such that $\frac{d-b}{2}\leq|F|\leq\frac{d-a}{2}$ and subsets of $[d-2|F|]$ such that $b\leq d-2|F|$ (equivalently $\frac{d-b}{2}\leq|F|$).
  2. 2.
    Now assume that $a\in V(S)$ and $b\in[d]$. The faces of $\Gamma$ affected are those using $a\in V(S)$ and $b\in[d]$. These come from faces $F\in S$ containing $a$ and subsets $G\in 2^{[d-2|F|]}$ containing $b$. In other words, we need $d-2|F|\geq b$ (equivalently $b\leq d-2|F|$ or $|F|\leq\frac{d-b}{2}$). For these faces, the subdividing vertex $v$ doubles the faces by replacing $a$ with $v$ or $b$ with $v$. The added part (faces of $\Gamma^{\prime}$ not counted by $\Gamma$) has something similar to a Boolean decomposition where we consider faces $F\in S$ that contain $a$ instead of all faces of $S$ and subsets of $[d-|F|]$ containing $b$ instead of all of $[d-2|F|]$.
  3. 3.
    If $a,b\in V(S)$, then this is essentially an edge subdivision $S^{\prime}$ of $S$ using the same $S$. The components/faces $G\in 2^{[d-2|F|]}$ used remain the same and we still have a Boolean decomposition without splitting the new simplicial complex $\Gamma^{\prime}$ into parts.
We now move to the other side and take $\Delta$ to be some flag Cohen–Macaulay simplicial complex. Write $h(\Delta)=f(\Gamma)$ for some (balanced) simplicial complex $\Gamma$ (which exists by work of Caviglia–Constantinescu–Varbaro [8] and Björner–Frankl–Stanley [6]). Given any flag simplicial complex $\Gamma$, there is a balanced flag vertex decomposable simplicial complex $\Delta$ such that $h(\Delta)=f(\Gamma)$ by a result of Constantinescu–Varbaro (Proposition 4.1 on p. 98 – 99 of [10]). Let $\Delta^{\prime}$ be the (stellar) edge subdivision of $\Delta$ at an edge $e\in\Delta$. Suppose that there are simplicial complexes $\Gamma$ and $\Gamma_{e}$ with Boolean decompositions
$\Gamma=\{F\cup G:F\in S,G\in 2^{[d-2|F|]}\}$
and
$\Gamma_{e}=\{F_{e}\cup G_{e}:F_{e}\in S_{e},G_{e}\in 2^{[d-2|F_{e}|-2]}\}$
involving simplicial complexes $S$ of dimension $\frac{d}{2}-1$ and $S_{e}$ of dimension $\frac{d}{2}-2$ such that $f(\Gamma)=h(\Delta)$ and $f(\Gamma_{e})=h(\operatorname{lk}_{\Delta}(e))$. This would mean that $\gamma_{i}(\Delta)=f_{i-1}(S)$ and $\gamma_{i}(\operatorname{lk}_{\Delta}(e))=f_{i-1}(S_{e})$. Note that $\gamma_{\Delta^{\prime}}(t)=\gamma_{\Delta}(t)+t\gamma_{\operatorname{lk}_{\Delta}(e)}(t)$. In other words, we have that $\gamma_{\Delta^{\prime}}(t)=f_{S}(t)+tf_{S_{e}}(t)$.
Since $\Delta$ is Cohen–Macaulay, we have that $h_{i}(\operatorname{lk}_{\Delta}(e))\leq h_{i}(\Delta)$ (Corollary 9.2 on p. 127 of [30]). This means that $f_{i-1}(S_{e})\leq f_{i-1}(S)$ and we can treat $S_{e}$ as a subcomplex of $S$ by using compression complexes. Looking at the construction for simplicial complexes whose $f$-vectors yield $h$-vectors of flag vertex decomposable simplicial complexes from Theorem 3.3 on p. 96 – 97 of [10], we see that this implies the existence of a Boolean decomposition for a simplicial complex with the same $f$-vector as $\Gamma^{\prime}$ by setting $S^{\prime}=S\cup(S_{e}*u)$ for some disjoint new vertex $u$. Going backwards from $\Gamma^{\prime}$ to $\Gamma$ is the same as removing $u$ (equivalently identifying $u$ with some vertex of $\Gamma_{e}$).
In summary, we have the following:
Proposition 1.6.
Let $\Gamma_{1}$ be a simplicial complex of dimension $D-1$ and $\Gamma_{2}\leq\Gamma_{1}$ be subcomplex of dimension $D-2$. Let $u$ be a new disjoint vertex. Note that we can instead take $\Gamma_{2}$ to be a simplicial complex such that $f_{i}(\Gamma_{2})\leq f_{i}(\Gamma_{1})$ for all $i$ since we can take compression complexes. If $\Gamma_{1}$ and $\Gamma_{2}$ have Boolean decompositions, the simplicial complex
$\Gamma\coloneq\Gamma_{1}*_{\Gamma_{2}}u=\Gamma_{1}\cup(\Gamma_{2}*u)$
also has a Boolean decomposition.
Proof.
Let
$\Gamma_{1}=\{F_{1}\cup G_{1}:F_{1}\in S_{1},G_{1}\in 2^{[d-2|F_{1}|]}\}$
and
$\Gamma_{2}=\{F_{2}\cup G_{2}:F_{2}\in S_{2},G_{2}\in 2^{[d-2|F_{2}|-1]}\}$
be Boolean decompositions for $\Gamma_{1}$ and $\Gamma_{2}$. Then, we have that $\Gamma_{2}*u$ has the Boolean decomposition given by
$\Gamma_{2}*u=\{F_{2}\cup\widetilde{G}_{2}:F_{2}\in S_{2},\widetilde{G}_{2}\in 2^{[d-2|F_{2}|]}\}.$
This yields the Boolean decomposition
$\Gamma=\{F\cup G:F_{1}\in S_{1}\cup S_{2},G\in 2^{[d-2|F|]}\}$
for $\Gamma$.
We can use this result and to apply the reasoning used in the edge subdivision case to geometrically natural examples. The statements below follow from the reasoning above and the proofs of Theorem 3.3 on p. 96 – 97 and Theorem 4.3 on p. 99 – 100 of [10].
Corollary 1.7.
  1. 1.
    Suppose that a Cohen–Macaulay simplicial complex $\Delta$ has an associated simplicial complex $\Gamma$ such that $h(\Delta)=f(\Gamma)$ and $\Gamma$ has a Boolean decomposition. Then, the same is true for an edge subdivision $\Delta^{\prime}$ of $\Delta$. In other words, there is a simplicial complex $\Gamma^{\prime}$ such that $h(\Delta^{\prime})=f(\Gamma^{\prime})$ and $\Gamma^{\prime}$ has a Boolean decomposition.
  2. 2.
    Suppose that $\Delta$ is a simplicial complex satisfying one of the following properties:
    • $\Delta$ is vertex decomposable and flag.
    • $\Delta$ is a balanced, vertex decomposable, flag simplicial complex.
    Any simplicial complexes associated to antistars and links of vertices with Boolean decompositions can be used to construct a flag (resp. quasi-flag from Definition 4.2 on p. 99 of [10]) simplicial complex $\Gamma$ such that $h(\Delta)=f(\Gamma)$ and $\Gamma$ has a Boolean decomposition.
2 Geometry via Tchebyshev transformations and subdivisions
Using the expression of the gamma polynomial as an inverted Chebyshev expansion and work of Hetyei [18] on Tchebyshev subdivisions, we can give geometric information on simplicial complexes $T$ such that $\gamma(\Delta)=f(T)$.
2.1 Background related to cell and cubical complexes
We give some background on objects that we will be studying. Recall that the $w=2\alpha+1$ substitution was motivated by trying to use $f_{S}(x)=F_{S}(2x+1)$ for a simplicial complex $S$ to show that the input for the Tchebyshev transformation (for polynomials) in $T(P_{\Delta}(w))=g_{\Delta}(2w)=(2w+2)^{\frac{d}{2}}\gamma_{\Delta}\left(\frac{1}{2w+2}\right)$ is the $F$-vector of some simplicial complex. However, the indices from Theorem 2.22 are flipped in that we have the index $\frac{d}{2}-k-1$ term for the coefficient of $\alpha^{k}$ instead of something of index $k-1$.
A cubical complex that is closely connected to $F$-polynomials we are considering the mirroring/Danzer complex $M\Delta$ of a cell complex $\Delta$, which is a cubical complex whose vertices have links isomorphic to $\Delta$.
Definition 2.1.
  1. 1.
    By cell complex, we mean a polyhedral complex or the (polar) dual of one (analogous to polar duals of polytopes).
  2. 2.
    We will take the dual $A^{*}$ of a cell complex $A$ of dimension $D-1$ to mean the cell complex where a face of dimension $j$ corresponds to the collection of facets of $A$ that contain a particular $(D-j-1)$-dimensional face of $A$. This is an analogue of the polar dual of a polytope.
We can also give a generalization of polar duality for multicomplexes.
Definition 2.2.
  1. 1.
    (Definition 2.4 on p. 25 – 26 of [6]) A multicomplex is a collection of monomials $\mathcal{M}$ such that $m|m^{\prime}\in\mathcal{M}$ implies that $m\in\mathcal{M}$ for all monomials $m^{\prime}\in\mathcal{M}$.
    Let $\mathcal{M}^{r}$ denote the set of all monomials in $\mathcal{M}$ of degree $r$. The $f$-vector of a multiplex $\mathcal{M}$ is the sequence $(f_{-1},f_{0},f_{1},\ldots)$ where $f_{j-1}(\mathcal{M})=|\mathcal{M}^{j}|$ for $j=0,1,2,\ldots$. Note that this uses a different indexing convention from [6].
  2. 2.
    The dual $\mathcal{M}^{*}$ of a degree $D$ multicomplex $\mathcal{M}$ is the complex where the 0-dimensional faces/vertices come from maximal degree monomials and the $(k-1)$-dimensional faces of $\mathcal{M}^{*}$ come from quotients of degree $D$ terms of $\mathcal{M}$ by degree $D-k$ terms of $\mathcal{M}$ In other words, the“vertices” contained in a face $R^{*}\in\mathcal{M}$ of degree $k$ corresponding to a monomial $R\in\mathcal{M}$ come from to monomials of maximal degree in $\mathcal{M}$ that are divisible by $R$.
Definition 2.3.
(a small generalization of the definition in p. 299 of [5], p. 20 – 21 of [12], and p. 10 – 11 of [13])
Let $T$ be a cell complex. We can think of $T$ (including the empty set) as a partially ordered set (the corresponding simplicial poset) where each face is a 0-1 vector with a 0 in the $i^{\text{th}}$ place if and only if the vertex $i$ belongs to the face. Thus $T$ is partially ordered by $0>1$ extended componentwise.
Then we construct a partially ordered set $MT$ (called the mirroring, Danzer complex, or power complex $2^{T}$ of $T$) as follows:
$MT\coloneq\{(a_{1},\ldots,a_{n}):(|a_{1}|,\ldots,|a_{n}|)\in T\}\subset I^{n},$
partially ordered by $0>1$ and $0>-1$ extended componentwise. Here $I$ is the poset $\{0,1,-1\}$ with ordering $1<0$ and $-1<0$. Note that $\dim MT=\dim T+1$.
Remark 2.4.
The way we will think about the notation above is that placing $1$ or $-1$ in the $i^{\text{th}}$ coordinate refers to intersecting with the hyperplane $x_{i}=1$ or $x_{i}=-1$ defining a facet of the $n$-cube $[-1,1]^{n}$ and that the coordinates labeled with a “0” are “free” in the sense that they are used to span a face of the cubical complex $MT$. In Rowlands’ thesis (see p. 4 of [28]), the “free” coordinates are labeled with a “*”, which may better convey this. As a consequence, the faces of $MT$ with 00’s in the same coordinates (i.e. coming from the same face of $T$) are parallel to each other.
Example 2.5.
(Intersection patterns of faces on $M\Delta$)
  1. 1.
    Let $[n]$ be the vertex set of $\Delta$. A collection of faces $G_{1},\ldots,G_{m}\in M\Delta$ has a nonempty intersection if and only if the following condition holds: For all $1\leq i\leq n$, $(G_{k})_{i},(G_{\ell})_{i}\neq 0\Longrightarrow(G_{k})_{i}=(G_{\ell})_{i}$.
  2. 2.
    Suppose that the faces $G_{1},\ldots,G_{m}\in MT$ have a nonempty intersection. For each $1\leq i\leq m$, let $F_{i}=\{(|a_{1}|,\ldots,|a_{n}|):(a_{1},\ldots,a_{n})\in G_{i}\}$ (i.e. subset labeled by 0’s corresponds to $F_{i}$). Then, we have that $G_{1}\cap\cdots\cap G_{m}$ has 0’s in the same slots as $F_{1}\cap\cdots\cap F_{m}\in\Delta$ and the “common values” from Part 1 for coordinates corresponding to elements of the complement $[n]\setminus(F_{1}\cap\cdots\cap F_{m})$. In particular, this means that $G_{1}\cap\cdots\cap G_{m}\in M\Delta$ since the subset of $[n]$ labeled with 0’s is the face $F_{1}\cap\cdots\cap F_{m}\in\Delta$ and the rest of the coordinates are labeled with ±$\pm 1$.
  3. 3.
    In inclusions of faces of $M\Delta$, the changes in coordinates can only come from additional 0’s being placed that are induced by inclusions of faces of $\Delta$. Given a face $G\in M\Delta$. This implies the following:
    • $\operatorname{st}_{M\Delta}(G)\cong\operatorname{st}_{\Delta}(|G|)$, where $|G|$ denotes the face of $\Delta$ obtained after taking the absolute value of each coordinate (i.e. setting the $-1$ coordinates of $G$ equal to 1).
    • The nonzero coordinates of elements of $\operatorname{st}_{M\Delta}(G)$ are determined by those of $G$ (which are indexed by $[n]\setminus|G|$).
Returning to the original question of the interpretation of $P_{\Delta}(2\alpha+1)$, we use the following result connecting faces of a cell complex with its mirroring/Danzer complex.
Definition 2.6.
($\widetilde{f}$-polynomial of a poset, p. 307 of [5])
Given a poset $P$ such that $\operatorname{rank}(P)=D$, we denote by $\widetilde{f}_{i}\coloneq\widetilde{f}_{i}(P)$ the number of elements of rank $i$ and define the polynomial
$\widetilde{f}(P,t)\coloneq\sum_{i=0}^{D}\widetilde{f}_{i}t^{i}.$
Remark 2.7.
(Comments on indexing conventions for simplicial vs. cubical complexes)
We write $\widetilde{f}_{i}$ in a change from the notation in the original work of Babson–Billera–Chan [5] since there may be a difference from the usual definition of the $f$-polynomial of a simplicial complex due to the indexing conventions used there for face posets of simplicial complexes and cubical complexes (see discussion on p. 298 of [5]). To be more precise, the empty set is included in the face poset of a simplicial complex while it is not included in the face poset of a cubical complex. This is to preserve the property of the order ideals of these face posets being Boolean algebras and products of copies of the face poset of an interval respectively. Another source where the poset $\widetilde{f}$-polynomial is used is on p. 13 of work of Rowlands [29] on connections between $CAT(0)$ cubical complexes and flag simplicial complexes (where $f$ is taken to mean $\widetilde{f}$ in our setting).
As a result, we have that $\widetilde{f}_{i}(S)=f_{i-1}(S)$ for a simplicial complex $S$ and $\widetilde{f}_{j}(K)=f_{j}(K)$ for a cubical complex $K$ when we take $f_{\ell}(A)$ to mean the number of $\ell$-dimensional faces of a polyhedral complex $A$ as usual. We take a look at what this means for the comparison between
$\widetilde{f}(P,t)\coloneq\sum_{i=0}^{D}\widetilde{f}_{i}(P)t^{i}$
and the “usual” $f$-polynomial
$f_{A}(t)\coloneq\sum_{i=0}^{D}f_{i-1}(A)t^{i}.$
Since $\widetilde{f}_{i}(S)=f_{i-1}(S)$ for a simplicial complex $S$, we actually have that $\widetilde{f}(S,t)=f_{S}(t)$. However, we need to make a modification for a cubical complex $K$ since $\widetilde{f}_{i}(K)=f_{i}(K)$ for a cubical complex $K$. In that case, we have that
$\widetilde{f}(K,t)$ $=\sum_{i=0}^{D}\widetilde{f}_{i}(K)t^{i}$
$=\sum_{i=0}^{D}f_{i}(K)t^{i}$
$=f_{0}(K)+f_{1}(K)t+f_{2}(K)t^{2}+\ldots+f_{D}(K)t^{D}$
$\Longrightarrow t\widetilde{f}(K,t)$ $=f_{0}(K)t+f_{1}(K)t^{2}+f_{2}(K)t^{3}+\ldots+f_{D}(K)t^{D+1}$
$=f_{D}(K)-1$
$\Longrightarrow f_{K}(t)$ $=t\widetilde{f}(K,t)+1.$
Using this notation, there is a clean way to connect the $\widetilde{f}$-polynomials of a cell complex $T$ and its mirroring/Danzer complex $MT$.
Proposition 2.8.
(extension of p. 302 of [5])
Suppose that we include the empty set in the face poset of a cell complex $T$ and that it is not included in the face poset of $MT$. If every $(k-1)$-dimensional cell of $T$ has $k$ vertices, we have that
$\widetilde{f}(MT,t)=2^{n(T)}\widetilde{f}\left(T,\frac{t}{2}\right),$
the notation of Definition 2.6 where $n(T)$ denotes the number of vertices in $T$.
In general, we have that
$\widetilde{f}(MT,t)=2^{n(T)}\widetilde{S}\left(T,\frac{t}{2}\right),$
where $\widetilde{S}(T,x)=\sum_{i=0}^{n(T)}\widetilde{s}_{i}x^{i}$ with $\widetilde{s}_{i}\coloneq\#\{F\in T:|F|=i\}$.
2.2 Colorings, Tchebyshev subdivisions, and intervals
Now suppose that $\gamma(\Delta)=f(T)$ for some simplicial complex $T$ of dimension $\frac{d}{2}-1$. By Theorem 0.3, we have that
$g_{\Delta}(2w)=(2w+2)^{\frac{d}{2}}\gamma_{\Delta}\left(\frac{1}{2w+2}\right)=(2w+2)^{\frac{d}{2}}f_{T}\left(\frac{1}{2w+2}\right).$
$g_{\Delta}(2w)$ $=\sum_{i=0}^{\frac{d}{2}}f_{i-1}(T)(2w+2)^{\frac{d}{2}-i}$
$=\sum_{i=0}^{\frac{d}{2}}f_{i-1}(T)2^{\frac{d}{2}-i}(w+1)^{\frac{d}{2}-i}$
$=\sum_{i=0}^{\frac{d}{2}}\widetilde{f}_{i}(T)2^{\frac{d}{2}-i}(w+1)^{\frac{d}{2}-i}$
$=\sum_{j=0}^{\frac{d}{2}}\widetilde{f}_{\frac{d}{2}-j}(T)2^{j}(w+1)^{j}$
$\Longrightarrow[w^{k}]g_{\Delta}(2w)$ $=\sum_{j=k}^{\frac{d}{2}}\widetilde{f}_{\frac{d}{2}-j}(T)2^{j}\binom{j}{k}$
$=\sum_{j=k}^{\frac{d}{2}}\widetilde{f}_{j}(T^{*})2^{j}\binom{j}{k}$
since $\widetilde{f}_{i}(S)=f_{i-1}(S)$ for a simplicial complex $S$ by Remark 2.7.
We note that this has some resemblance to first cubical barycentric subdivision of a simplicial complex mentioned defined by Hetyei (see [16], which is defined below). We will also give a generalization of Babson–Billera–Chan [5] which puts it in context of well-known subdivisions.
Definition 2.9.
(Hetyei, Définition 6 on p. 45 and Remarque 5 on p. 45 – 46 of [16])
The first cubical barycentric subdivision of a simplicial complex $S$ of dimension $D-1$ is the cubical complex $\mathcal{C}(S)$ defined as follows:
  1. 1.
    The set of vertices of $\mathcal{C}(S)$ is the set of nonempty faces of $S$.
  2. 2.
    The faces of $\mathcal{C}(S)$ are the intervals of the face poset of $S$, that is all sets of the form
    $[\sigma,\tau]\coloneq\{\lambda\in S:\sigma\subset\lambda\subset\tau\}.$
We can express the $f$-vector $(f_{-1}^{\Box}(S),\ldots,f_{D-1}^{\Box}(S))$ of the complex $\mathcal{C}(S)$ in terms of the $f$-vector $(f_{-1}(S),\ldots,f_{d-1}(S))$ of $S$ as follows:
$f_{k}^{\Box}(S)=\sum_{j=k}^{d-1}\binom{j+1}{k}f_{j}(S)$
Remark 2.10.
(Comparison with the formula for $[w^{k}]g(2w)$ given if $\gamma(\Delta)=f(T)$)
  1. 1.
    Comparing this to the formula for $[w^{k}]g(2w)$ obtained when $\Gamma(\Delta)=f(T)$, we see that it uses $\binom{j}{k}$ instead of $\binom{j+1}{k}$ and has an input that might not be a simplicial complex. Note that the $k$-dimensional faces of $MA$ come from $(k-1)$-dimensional faces of $A$ (see Definition 2.3).
  2. 2.
    The cubical barycentric subdivision assumes that both endpoints of the interval considered are contained in $S$. If we drop the condition that $S$ is a simplicial complex, we’re simply considering endpoints $[\sigma,\tau]$ where the upper endpoint is a $j$-dimensional face of $S$ and $\tau$ is some subset of $\sigma$ (which may or may not be a face of $S$).
A generalization by Babson–Billera–Chan [5] is defined below.
Definition 2.11.
(Babson–Billera–Chan, Section 2.3 on p. 303 – 304 of [5])
Given a poset $P$, its barycentric cover $KP$ is the poset with elements the order relations of $P$, partially ordered by inclusion (i.e. $(u\leq v)\leq(x\leq y)$ (see Definition 2.14) if and only if $x\leq u\leq v\leq y$). Note that $KP=K^{\text{op}}$ (see p. 298 of [5]) and $K(P\times Q)=K(P)\times K(Q)$.
Here are some properties of the barycentric cover $KP$.
Proposition 2.12.
(Babson–Billera–Chan, Proposition 2.2 on p. 303 of [5])
If $P$ has Boolean intervals, then $KP$ is a cubical poset. This includes simplicial and cubical posets (more generally, face posets of polyhedral complexes with simple, nonempty cells) as well as their duals. Further, if $\widehat{P}$ ($P$ with $\widehat{0}$ and $\widehat{1}$ adjoined) is a lattice, then so is $\widehat{KP}$.
Theorem 2.13.
(Babson–Billera–Chan, Theorem 2.3 and Corollary 2.4 on p. 303, p. 298 of [5])
  1. 1.
    Given a poset $Q$, let $|Q|$ be the (simplicial) complex of chains in $P$ (i.e. the order complex). Then, $|KP|$ gives a polyhedral subdivision of $|P|$ by the map taking the point $(x\leq x)$ to itself and the point $(x<y)$ in $|KP|$ to the midpoint of the edge $x<y$ in $|P|$, and extending linearly over every closed simplex in $|KP|$.
  2. 2.
    If $P$ is the face poset of a polyhedral complex, then $KP$ is the face poset of a polyhedral subdivision of $P$, lying between $P$ and $|P|$ in the refinement order.
These interval constructions are closely connected to Tchebyshev subdivisions by recent work of Hetyei [19].
Definition 2.14.
(Definition 8 on p. 924 of [19])
An interval $[u,v]$ in a partially ordered set $P$ is the set of all elements $w\in P$ satisfying $u\leq v\leq w$. For a finite partially ordered set $P$, we define the poset $I(P)$ of the nonempty intervals of $P$ as the set of all intervals $[u,v]\subset P$, ordered by inclusion.
The connection to Tchebyshev subdivisions comes from taking order complexes.
Theorem 2.15.
(Walker–Jojić–Hetyei, Theorem 9 on p. 925 of [19])
For any finite partially ordered set $P$, the order complex $\Delta(I(P))$ of its poset of nonempty intervals is isomorphic to a Tchebyshev triangulation of $\Delta(P)$ as follows:
For each $u\in P$, we identify the vertex $[u,u]\in\Delta(I(P))$ with the vertex $u\in\Delta(P)$ and for each nonsingleton interval $[u,v]\in I(P)$, we identify the vertex $[u,v]\in\Delta(I(P))$ with the midpoint of the edge $\{[u,u],[v,v]\}$. We number the midpoints $[u_{1},v_{1}],[u_{2},v_{2}],\ldots$ in such a way that $i<j$ holds whenever the interval $[u_{i},v_{i}]$ contains the interval $[u_{j},v_{j}]$.
We have a similar situation for intervals of graded posets.
Definition 2.16.
(Definition 10 on p. 926 of [19])
For a graded poset $P$ we define its poset of intervals $\widehat{I}(P)$ as the poset of all intervals of $P$, including the empty set, ordered by inclusion.
Proposition 2.17.
(Hetyei, Proposition 12 on p. 927 of [19])
Let $P$ be a graded poset and $\widehat{I}(P)$ be its poset of intervals. Then the order complex $\Delta(\widehat{I}(P)-\{\emptyset,[\widehat{0},\widehat{1}]\})$ is a Tchebyshev triangulation of the suspension of $\Delta(P-\{\widehat{0},\widehat{1}\})$.
Remark 2.18.
(Structure of $g_{\Delta}(2w)$ expression after substituting in $f(T)$ for some simplicial complex $T$)
We take a page from p. 13 of Rowlands’ paper [29], which shows that $\widetilde{f}(\Box_{P},t)=\widetilde{f}(\Delta_{P},t+1)$. Note that any flag complex is the crossing complex $\Delta_{P}$ of a $PIP$ $P$ (Lemma 3.3 on p. 11 of [29]) and any rooted $CAT(0)$ cubical complex is of the form $\Box_{P}$ for some $P$ (Theorem 2.9 on p. 9 of [29]). The proof uses a bijection between dimension $i$ elements of $\Box_{P}$ and pairs $(A,M)$, where $A$ is the set of consistent antichains of $P$ and $M\subset A$ is a subset with $|M|=i$. The identity follows after making a change in the order of summation.
Another way to say this is that
$[w^{k}]\widetilde{f}\left(T(A),\frac{w-1}{2}\right)=\sum_{j=k}^{\frac{d}{2}}\widetilde{f}_{j}(T^{*})2^{j}\binom{j}{k},$
where $A$ is cell complex related to unused colors of faces in a balanced simplicial complex with an assigned coloring. Note that the last expression makes sense since a $(k-1)$-face of $T(A)$ has $k$ vertices (see Definition 2.26).
We can apply the setup above to give an interpretation of $P_{\Delta}(2\alpha+1)$, where
$P_{\Delta}(u)\coloneq h_{\frac{d}{2}}(\Delta)+2\sum_{j=1}^{\frac{d}{2}}h_{\frac{d}{2}-j}(\Delta)u^{j}.$
The motivation is to show that $P_{\Delta}(u)$ is the $F$-polynomial of some cell complex $A$. Since $f_{S}(x)=F_{S}(2x+1)$, this is equivalent to showing that $P_{\Delta}(2\alpha+1)$ is the $f$-polynomial of some cell complex $A$. As it turns out, the Tchebyshev subdivisions studied by Hetyei [18] can be generalized to arbitrary cell complexes (in the sense of Definition 2.1). This implies that $g(2w)=F_{T(A)}(w)$. Note that $g(2w)=(2w+2)^{\frac{d}{2}}\gamma\left(\frac{1}{2w+2}\right)$.
Before moving on, we first record a construction that is used to express $P$ in terms of a modification of the $f$-vector of some cell complex.
Definition 2.19.
(Signed unused color complex)
Let $\Gamma$ be a $(D-1)$-dimensional balanced simplicial complex. Fix a $D$-coloring of the $1$-skeleton of $\Gamma$. We define the signed unused color complex $\mathcal{D}(\Gamma)$ of $\Gamma$ to be the cell complex below.
  • The vertices are of the form $\{q\}\subset\{q\}\subset C_{F}$ or $\emptyset\subset\{q\}\subset C_{F}$, where $C_{F}\subset[D]$ is the set of colors unused by $F$. In the first case, we will denote a vertex by $v^{+}$ and will write $v^{-}$ in the second case. Note that $C_{\emptyset}=[D]$ since none of the colors are used.
  • The $(k-1)$-faces of the form $B\subset Q\subset C_{F}$, where $C_{F}$ is the set of colors in $[D]$ not used by $F$, $Q\subset C_{F}$ is a $k$-element subset, and $B\subset Q$ is an arbitrary subset. This is like considering “signed” $k$-element subsets of $C_{F}$. Note that $C_{F}\neq\emptyset$ since $Q$ is nonempty for the cases we will consider.
  • We have an inclusion of faces $(B_{1}\subset Q_{1}\subset C_{F_{1}})\leq(B_{2}\subset Q_{2}\subset C_{F_{2}})$ if and only if $F_{1}\supset F_{2}$, $B_{1}\subset B_{2}$, and $Q_{1}\setminus B_{1}\setminus Q_{2}\setminus B_{2}$.
Next, we describe when a collection of vertices of $\mathcal{D}(\Gamma)$ forms a face of $\mathcal{D}(\Gamma)$.
Proposition 2.20.
Let $\Gamma$ be a balanced simplicial complex of dimension $D-1$. A collection of vertices $v_{1}^{\varepsilon_{1}},\ldots,v_{k}^{\varepsilon_{k}}\in V(\mathcal{D}(\Gamma))$ with $\varepsilon_{i}=\pm 1$ forms a face of $\mathcal{D}(\Gamma)$ if and only if one of the following conditions (in the notation of Definition 2.19) holds:
  1. 1.
    $F_{1}\cap\cdots\cap F_{k}\neq\emptyset$.
  2. 2.
    $F_{1}\cap\cdots\cap F_{k}=\emptyset$ and $C_{i_{1}}\cup\cdots\cup C_{i_{k}}=[D]$.
Note that each $C_{i}$ is determined by $F_{i}\in\Gamma$.
Remark 2.21.
The second condition is nontrivial. For example, suppose that $\Gamma$ is the boundary of the octahedron/2-cross polytope (meaning $D=3$). This is a balanced simplicial complex. Color the vertices ±$\pm e_{i}$ with $i$ for $1\leq i\leq 3$ by $i$ to get a 3-coloring of its 1-skeleton. Consider the vertex $F_{1}=\{-e_{2}\}$ and the edge $F_{2}=\{e_{2},e_{3}\}$. Then, we have that $C_{F_{1}}=\{1,3\}$ and $C_{F_{2}}=\{1\}$. Since $F_{1}\cap F_{2}=\emptyset$ and $C_{F_{1}}\cup C_{F_{2}}=\{1,3\}$, the union of any two vertices of $\mathcal{D}(\Gamma)$ associated to these faces cannot form a face of $\mathcal{D}(\Gamma)$.
Proof.
Each vertex is of the form $B_{i}\subset\{q_{i}\}\subset C_{F_{i}}$, where $C_{F_{i}}$ is the set of unused colors of $F_{i}\in\Gamma$, $q_{i}\in C_{F_{i}}$, and $B_{i}=\{q_{i}\}$ or $\emptyset$. Using the inclusion relation in Definition 2.19, we have that the union $\{v_{1}^{\varepsilon_{1}},\ldots,v_{k}^{\varepsilon_{k}}\}$ is given by $B_{1}\cup\cdots\cup B_{k}\subset\{q_{1},\ldots,q_{k}\}\subset(C_{1}\cup\cdots\cup C_{k})_{F_{1}\cap\cdots\cap F_{k}}$. If $F_{1}\cap\cdots\cap F_{k}\neq\emptyset$, this forms a face of $\mathcal{D}(\Gamma)$ since $(P\cap Q)^{c}=P^{c}\cup Q^{c}$ for $P,Q\subset[D]$. Now suppose that $F_{1}\cap\cdots\cap F_{k}=\emptyset$. As mentioned in Definition 2.19, we have that $C_{\emptyset}=[D]$. This is only possible if $C_{F_{1}}\cup C_{F_{k}}=[D]$.
We can use this cell complex to connect $P_{\Delta}(u)$ to the $f$-vector of some cell complex.
Theorem 2.22.
Let $\Delta$ be a flag simplicial homology sphere of even dimension $d$ and
$P_{\Delta}(u)\coloneq h_{\frac{d}{2}}(\Delta)+2\sum_{j=1}^{\frac{d}{2}}h_{\frac{d}{2}-j}(\Delta)u^{j},$
which is $g(2u)$ with $T_{k}(u)$ replaced by $u^{k}$ (i.e. $g(2u)=T(P(u))$ and $P(u)=T^{-1}g(2u)$).
Given $k\geq 1$, there is a $\left(\frac{d}{2}-1\right)$-dimensional balanced simplicial complex $\Gamma$ such that
$[\alpha^{k}]\frac{P_{\Delta}(2\alpha+1)+f_{\frac{d}{2}-1}(\Gamma)}{2}-\sum_{j=0}^{\frac{d}{2}}f_{\frac{d}{2}-j-1}(\Gamma)+1=f_{k-1}(\mathcal{D}(\Gamma))$
and
$\frac{P_{\Delta}(2\alpha+1)+f_{\frac{d}{2}-1}(\Gamma)}{2}-\sum_{j=0}^{\frac{d}{2}}f_{\frac{d}{2}-j-1}(\Gamma)+1=f_{\mathcal{D}(\Gamma)}(\alpha).$
Substituting in $\alpha=\frac{\beta-1}{2}$, we have that
$\frac{P_{\Delta}(\beta)+f_{\frac{d}{2}-1}(\Gamma)}{2}-\sum_{j=0}^{\frac{d}{2}}f_{\frac{d}{2}-j-1}(\Gamma)+1=F_{\mathcal{D}(\Gamma)}(\beta),$
where
$F(u)\coloneq\sum_{i=0}^{D}f_{i-1}(A)\left(\frac{u-1}{2}\right)^{i}$
for a cell complex $A$ of dimension $D-1$.
Proof.
Since $\Delta$ is a flag simplicial homology sphere, it is also a flag Cohen–Macaulay simplicial complex (e.g. see p. 21 of [4]). Then, a combining a result of Caviglia–Constantinescu–Varbaro (Corollary 2.3 on p. 474 of [8]) with that of Björner–Frankl–Stanley ($a=(1,\ldots,1)$ case of Theorem 1 on p. 23 and p. 33 of [6]) implies that $h(\Delta)$ is the $f$-vector of a balanced $\left(\frac{d}{2}-1\right)$-dimensional simplicial complex $\Gamma$.
Substituting in $x=2\alpha+1$, we have
$P_{\Delta}(2\alpha+1)$ $=\sum_{i=0}^{\frac{d}{2}}f_{i-1}(\Gamma)(2\alpha+1)^{\frac{d}{2}-i}$
$=\sum_{j=0}^{\frac{d}{2}}f_{\frac{d}{2}-j-1}(\Gamma)(2\alpha+1)^{j}.$
This implies that
$[\alpha^{k}]P_{\Delta}(2\alpha+1)$ $=\sum_{j=k}^{\frac{d}{2}}f_{\frac{d}{2}-j-1}(\Gamma)2^{k}\binom{j}{k}$
$=2^{k}\sum_{j=k}^{\frac{d}{2}}f_{\frac{d}{2}-j-1}(\Gamma)\binom{j}{k}.$
We claim that the $f$-vector of the complex $\mathcal{D}(\Gamma)$ from Definition 2.19 has dimensions given by the sum above. Given a $\left(\frac{d}{2}-j-1\right)$-dimensional $F\in\Gamma$, there are $j$ unused colors forming the set $C_{F}\subset\left[\frac{d}{2}\right]$. Choosing $k$ of them gives the subset $Q\subset C_{F}$ from Definition 2.19 and is counted by the term $f_{\frac{d}{2}-j-1}(\Gamma)\binom{j}{k}$ in the sum. Assigning a $\pm$ signs to each element of $Q$ gives the multiplication by $2^{k}$.
We record the (extended) $F$-polynomial definition to give some geometric context.
Definition 2.23.
(Definition 3.2 on p. 578) Given a cell complex (Definition 2.1) $A$ of dimension $D-1$, the $F$-polynomial of $A$ is
$F_{A}(x)\coloneq\sum_{j=0}^{D}f_{j-1}(A)\left(\frac{x-1}{2}\right)^{j}.$
Note that the $\widetilde{f}$-polynomial of the mirroring/Danzer complex $MA$ of $A$ has a close connection to the $F$-polynomial of $A$.
Proposition 2.24.
Given a cell complex $A$ such that $(k-1)$-dimensional faces have $k$ vertices, we have that
$F_{A}(x)=\frac{1}{2^{n(A)}}\widetilde{f}(MA,x-1),$
where $MA$ is the mirroring/Danzer complex associated to $S$ (Definition 2.3).
Equivalently, we have that
$\widetilde{f}(MA,x)=2^{n(A)}F_{A}(x+1).$
Since $f_{K}(t)=t\widetilde{f}(K,t)+1$ for a cubical complex $K$ by the discussion in Remark 2.7, this implies that
$f(MA,x)=2^{n(A)}xF_{A}(x+1)+1$
for such a cell complex $A$.
Remark 2.25.
We can write a generalization applying to arbitary cell complexes using Proposition 2.8 if we replace the $\widetilde{f}$-polynomial with the $\widetilde{S}$-polynomial parametrized by the number of faces with a given number of vertices (instead of dimension).
Proof.
In spite of the indexing issues discussed in Remark 2.7, the $\widetilde{f}$-polynomial of a cell complex from Definition 2.6 is the same as its usual $f$-polynomial. Now recall from Proposition 2.8 that
$\widetilde{f}(MT,t)=2^{n}\widetilde{f}\left(T,\frac{t}{2}\right)$
for a cell complex $T$ with $n$ vertices such that $(k-1)$-dimensional faces have $k$ vertices.
Since $F_{A}(x)=f_{A}\left(\frac{x-1}{2}\right)$ for a cell complex $A$, this means that
$\widetilde{f}(MA,t)$ $=2^{n}\widetilde{f}\left(A,\frac{t}{2}\right)$
$=2^{n}f\left(A,\frac{t}{2}\right)$
$\Longrightarrow\widetilde{f}(MA,x-1)$ $=2^{n}f\left(A,\frac{x-1}{2}\right)$
$=2^{n}F_{A}(x)$
$\Longrightarrow F_{A}(x)$ $=\frac{1}{2^{n}}\widetilde{f}(MA,x-1)$
after substituting in $t=x-1$.
The main application of the $F$-polynomial of a simplicial complex $\Delta$ in [18] is its compatibility with Tchebyshev transformations $T:\mathbb{R}[x]\longrightarrow\mathbb{R}[x]$ sending $x^{n}\mapsto T_{n}(x)$ (the $n^{\text{th}}$ Chebyshev polynomial of the first kind) since $T(F(\Delta))(x)=F_{T(\Delta)}(x)$ (Proposition 3.3 on p. 578 of [18]). As it turns out, the definition of a Tchebyshev subdivision and the formula for its $f$-polynomial (in the usual sense) do not depend on $\Delta$ being a simplicial complex. We will list the generalizations below and comment on why this is the case for the relevant definition and result.
Definition 2.26.
(Cell complex/multicomplex generalization of Definition 2.1 on p. 575 of [18])
Let $A$ be a cell complex from Definition 2.1 or the (dual of) a multicomplex from Definition 2.2.
Given a cell complex $A$ on a vertex set $V$, fix a linear order $<$ on $V$ and introduce a new element $\widehat{1}$ larger than all elements of $V$. The Tchebyshev triangulation $T_{<}A$ of $A$ is the cell complex on the vertex set
$\{(u,v):u,v\in V,u<v,\{u,v\}\subset F\text{ for some }F\in A\}\cup\{(u,\widehat{1}):u\in V\}$
whose $(k-1)$-faces are all sets $\{(u_{1},v_{1}),\ldots,(u_{k},v_{k})\}$ satisfying $v_{1}<\cdots<v_{k}$ and the following conditions:
  1. 1.
    The set $\{u_{1},v_{1},u_{2},v_{2},\ldots,u_{k},v_{k}\}\setminus\{\widehat{1}\}$ is a face of $A$.
  2. 2.
    For all $i\leq k-1$, either $u_{i}=u_{i+1}$ or $v_{i}\leq u_{i+1}$ holds.
Since the $f$-vector of $T_{<}(A)$ is independent of the ordering $<$, we will often drop the ordering and simply write $T(A)$.
Remark 2.27.
(Connection between definitions and comments on generalization to cell complexes)
  1. 1.
    This is an adaptation of Definition 2.1 on p. 575 of [18] by Hetyei which specializes to the original definition when $A$ is a simplicial complex. We can think of pairs of actual vertices of $A$ as vertices subdividing the corresponding edges and the pairs ending with $\widehat{1}$ correspond to vertices of the original cell complex $A$. The remaining conditions ensure that the “realization” lies inside some face of $A$ and we’re moving forward with respect to an edge ordering. Note that we think about faces of the subdivision in terms of which “side” of each edge we’re on. This is used later in the $f$-vector count of Tchebyshev subdivisions and is reminiscent of how one thinks about faces in barycentric subdivisions.
    While a more intuitive definition matching the description above more directly is given in Definition 2 on p. 921 of [19], we use the one above since it generalizes to arbitary cell complexes. For example, 1-dimensional faces in the dual of a cell complex may have more than one vertex unless the starting cell complex gives rise to a pseudomanifold without boundary (in which case every ridge is contained in 2 facets in the original cell complex – see Definition 3.15 on p. 24 of [30]).
  2. 2.
    While each face $G$ of $T(A)$ of dimension $k-1$ has the “right” number of vertices to be a $(k-1)$-dimensional face of a simplicial complex, $T(A)$ itself isn’t necessarily a simplicial complex. This is because removing the underlying set of a vertex of $G$ (given by a singleton or a pair) from the face of $A$ given by the union of the underlying sets of the vertices of $G$ might not yield a face of $A$ if $A$ isn’t a simplicial complex.
  3. 3.
    We have that $\dim T(A)+1$ is the maximum number of vertices of a face of $A$. This is because a union of vertices of $A$ coming from $k$ vertices of the form $(u_{i},v_{i})$ has $m\in[k,2k]$ elements (e.g. see the proof of Theorem 3.1 on p. 577 – 578 of [18]). For example, Definition 2.19 implies that $\dim T(\mathcal{D}(\Gamma))=\frac{d}{2}-1$ since any face of $\mathcal{D}(\Gamma)$ has $\leq\frac{d}{2}$ vertices.
Theorem 2.28.
(Cell complex generalization of Theorem 3.1 on p. 577 and Proposition 3.3 on p. 578 – 579 of [18] by Hetyei)
Let $A$ be a cell complex (Definition 2.1).
  1. 1.
    (Theorem 3.1 on p. 577 – 578 of [18])
    For $k\geq 1$, the number of $(k-1)$-faces of of $T(A)$ is given by
    $f_{k-1}(T(A))=\sum_{j=k}^{2k}f_{j-1}(A)\cdot 2^{2k-j-1}\left(\binom{k}{2k-j}+\binom{k-1}{2k-j}\right).$
  2. 2.
    (Proposition 3.3 on p. 578 – 579 of [18])
    The Tchebyshev transform of the $F$-polynomial of $A$ is the $F$-polynomial of its Tchebyshev triangulation:
    $T(F_{A})(x)=F_{T(A)}(x).$
Remark 2.29.
(Comments on generalizations)
  1. 1.
    For Part 1, the proof of Theorem 3.1 on p. 577 – 578 of [18] doesn’t use anything about the property defining an abstract simplicial complex (subsets of a face of the given complex being a face). In other words, the only input comes from the ordering properties defining Tchebyshev subdivisions and pairs of vertices contained in a given face of $A$ from Definition 2.26. After starting out by selecting a face to use for the set
    $u_{1},v_{1},u_{2},v_{2},\ldots,u_{k},v_{k}\}\setminus\{\widehat{1}\},$
    the rest of the work has to do what terms to pair together and in what order the pairs should be listed. The main structure of the proof has to do with the smallest face of the original cell complex $A$ that “contains” the given face of the Tchebyshev subdivision $T(A)$ and keeping track of possible slots for $u_{i}$’s and $v_{j}$’s (labeled with $+$ and -- signs), and how many strings with a given set of signs satisfy condition b) in Part 1 of Definition 2.26. This is sort of similar to how we keep track of faces on a barycentric subdivision.
  2. 2.
    The identity in Part 2 is purely a consequence of the formula in Part 1.
Combining these generalizations with Theorem 2.22 implies the following:
Corollary 2.30.
Recall that
$g_{\Delta}(u)\coloneq h_{\frac{d}{2}}(\Delta)+2\sum_{j=1}^{\frac{d}{2}}h_{\frac{d}{2}-j}(\Delta)T_{j}\left(\frac{u}{2}\right).$
We have that
$\frac{g_{\Delta}(2w)+f_{\frac{d}{2}-1}(\Gamma)}{2}-\sum_{i=0}^{\frac{d}{2}}f_{\frac{d}{2}-j-1}(\Gamma)+1=F(T(\mathcal{D}(\Gamma)),w)=\frac{1}{2^{n(T(\mathcal{D}(\Gamma)))}}\widetilde{f}(M(T(\mathcal{D}(\Gamma)),w-1),$
where the last equality follows from Proposition 2.24.
We will now consider the Tchebyshev triangulation (Definition 2.26) of the signed unused color complex from Definition 2.19. The following is a reproduction of Proposition 2.20 which will be relevant to our analysis.
Proposition 2.31.
Let $\Gamma$ be a balanced simplicial complex of dimension $D-1$. A collection of vertices $v_{1}^{\varepsilon_{1}},\ldots,v_{k}^{\varepsilon_{k}}\in V(\mathcal{D}(\Gamma))$ with $\varepsilon_{i}=\pm 1$ forms a face of $\mathcal{D}(\Gamma)$ if and only if one of the following conditions (in the notation of Definition 2.19) holds:
  1. 1.
    $F_{1}\cap\cdots\cap F_{k}\neq\emptyset$.
  2. 2.
    $F_{1}\cap\cdots\cap F_{k}=\emptyset$ and $C_{i_{1}}\cup\cdots\cup C_{i_{k}}=[D]$.
Note that each $C_{i}$ is determined by $F_{i}\in\Gamma$.
Going back to the sums under consideration, recall from Corollary 2.30 that
$\frac{g_{\Delta}(2w)+f_{\frac{d}{2}-1}(\Gamma)}{2}-\sum_{i=0}^{\frac{d}{2}}f_{\frac{d}{2}-i-1}(\Gamma)+1=F(T(\mathcal{D}(\Gamma)),w)=f\left(T(\mathcal{D}(\Gamma)),\frac{w-1}{2}\right)=\frac{1}{2^{n(T(\mathcal{D}(\Gamma)))}}\widetilde{f}(M(T(\mathcal{D}(\Gamma))),w-1).$
The coefficient of the degree $k$ term in
$R(w)\coloneq f\left(T(\mathcal{D}(\Gamma)),\frac{w}{2}\right)$
is $f_{k-1}(T(\mathcal{D}(\Gamma)))2^{-k}$. A $k$-tuple of vertices of $T(\mathcal{D}(\Gamma))$ forming a face of $\mathcal{D}(\Gamma)$ has between $k$ and $2k$ choices of signs to make for each vertex of the Tchebyshev subdivision from possible sizes in a cell complex $A$ of the union of vertices of $A$ coming from $k$ vertices of of the Tchebyshev subdivision $T(A)$ (analogous to considerations in the proof of Theorem 3.1 on p. 577 – 578 of [18]). This bound comes from having $2$ sign choices for every “valid” singleton of the form $(u,\widehat{1})$ from Definition 2.26 and $4$ sign choices when we have a valid pair $(u,v)$. This means that we have $2^{m}$ sign choices for some $k\leq m\leq 2k$. Effectively, what this means is that we’re considering $k$-tuples consisting of unsigned pairs of the form $(u,\widehat{1})$ and pairs $(u,v)$ where only the second coordinate is signed. Equivalently, $u$ and $v$ must have the same sign. This is a sort of “complex of one-sided intervals” $\widetilde{T}(\mathcal{D}(\Gamma))$ (a modification of the Tchebyshev subdivision $T(\mathcal{D}(\Gamma))$ of $\mathcal{D}(\Gamma)$ which we define below (with relevant parts reproduced from Definition 2.26).
Definition 2.32.
A vertex $v=(B\subset\{q\}\subset C_{F})$ of $\mathcal{D}(\Gamma)$ has positive sign if $B=\{q\}$ and it has negative sign if $B=\emptyset$. We denote these cases using the notation $\operatorname{sgn}(v)=1$ and $\operatorname{sgn}(v)=-1$ respectively.
  • The vertex set of $\widetilde{T}(\mathcal{D}(\Gamma))$ is
    $\{(u,v)\in V(\mathcal{D}(\Gamma)):u<v,\{u,v\}\subset F\text{ for some }F\in\mathcal{D}(\Gamma)\text{ and }\operatorname{sgn}(u)=\operatorname{sgn}(v)\}\cup\{(u,\widehat{1}):u\in V(\mathcal{D}(\Gamma))\}.$
  • The $(k-1)$-faces of $\widetilde{T}(\mathcal{D}(\Gamma))$ are all the sets $\{(u_{1},v_{1}),\ldots,(u_{k},v_{k})\}$ satisfying $v_{1}<\cdots<v_{k}$ and the following conditions:
    1. 1.
      The set $\{u_{1},v_{1},u_{2},v_{2},\ldots,u_{k},v_{k}\}\setminus\{\widehat{1}\}$ is a face of $\mathcal{D}(\Gamma)$.
    2. 2.
      For all $i\leq k-1$, either $u_{i}=u_{i+1}$ or $v_{i}\leq u_{i+1}$ holds.
The expression in Remark 2.18 is
$g_{\Delta}(2w)=\sum_{j=0}^{\frac{d}{2}}f_{\frac{d}{2}-j-1}(T)2^{j}(w+1)^{j}=\sum_{j=0}^{\frac{d}{2}}f_{j-1}(T^{*})2^{j}(w+1)^{j}.$
In some sense, this indicates that the mirroring/Danzer complex of the Tchebyshev subdivision $T(\operatorname{lk}_{\Delta}(e_{k})\big{\backslash}\left(\left(\bigcup_{i=1}^{k-1}\operatorname{lk}_{\Delta}(e_{i})\right)\cap\operatorname{lk}_{\Delta}(e_{k})\right))$ of $\mathcal{D}(\Gamma)$ is the interval poset of the simplicial complex $T$ such that $\gamma(\Delta)=f(T)$.
This would be equivalent to having
$[w^{k}]g_{\Delta}(2w)=\sum_{j=k}^{\frac{d}{2}}f_{j-1}(T^{*})2^{j}\binom{j}{k}$
Note that the terms above also count signed intervals where the upper bound is assigned a sign and the lower one (from a subset of size $k$) is unsigned (equivalently inheriting signs from the top one).
So far, we have shown that
$[w^{k}]\frac{g_{\Delta}(2w)+f_{\frac{d}{2}-1}(\Gamma)}{2}-\sum_{i=0}^{\frac{d}{2}}f_{\frac{d}{2}-i-1}(\Gamma)+1=F(T(\mathcal{D}(\Gamma)),w)=\sum_{j=k}^{\frac{d}{2}}f_{j-1}(A)2^{j}\binom{j}{k},$
where $A$ denotes the $A(w)$ such that $R(w)=A(w+1)$. We can take $A(w)$ to have nonnegative coefficients when $\gamma(\Delta)=f(T)$.
After multiplying by 2, we have
$[w^{k}]g_{\Delta}(2w)+f_{\frac{d}{2}-1}(\Gamma)-2\sum_{i=0}^{\frac{d}{2}-1}f_{\frac{d}{2}-i-1}(\Gamma)=2F(T(\mathcal{D}(\Gamma)),w)$ $=2\sum_{j=k}^{\frac{d}{2}}f_{j-1}(A)2^{j}\binom{j}{k}$
$\Longrightarrow[w^{k}]g_{\Delta}(2w)-\left(f_{\frac{d}{2}-1}(\Gamma)+2\sum_{i=1}^{\frac{d}{2}-1}f_{\frac{d}{2}-i-1}(\Gamma)\right)$ $=2\sum_{j=k}^{\frac{d}{2}}f_{j-1}(A)2^{j}\binom{j}{k},$
which means that we put $\pm$ signs on each face (alternatively put signs on an extraneous element $\widehat{1}$ for each face).
To obtain a cell complex $T^{*}$ where
$[w^{0}]g_{\Delta}(2w)=2\sum_{j=0}^{\frac{d}{2}}f_{j-1}(A)+\left(f_{\frac{d}{2}-1}(\Gamma)+2\sum_{i=1}^{\frac{d}{2}-1}f_{\frac{d}{2}-i-1}(\Gamma)\right)=\sum_{j=0}^{\frac{d}{2}}f_{j-1}(T^{*}),$
it suffices to formally “add” $\left(f_{0}(\Gamma^{*})+2\sum_{i=1}^{\frac{d}{2}-1}f_{i}(\Gamma^{*})\right)$ disjoint points since $f_{0}(T^{*})$ does not appear in the sums determining $[w^{k}]g_{\Delta}(2w)$ for any $k\geq 1$.
Corollary 2.33.
Under the modifications above, the mirroring/Danzer complex $M(T(\mathcal{D}(\Gamma)))$ is the (signed) interval complex/generalized cubical barycentric subdivision associated to the (dual of) a simplicial complex $S$ such that $\gamma(\Delta)=f(S)$ if one exists.
Finally, we study some geometric properties connected to nonnegativity of translation by $-1$ (thinking about the gamma vector). Given $MA$, note that the links of the vertices are isomorphic to $A$. The following result shows that $MA$ is a $CAT(0)$ cubical complex if and only if $A$ is a flag simplicial complex and $MA$ is simply connected:
Theorem 2.34.
(Gromov, Section 4 of [15] and Theorem 2.3 of [29])
A cubical complex $\Box$ is $CAT(0)$ if and only if it is simply connected and all the vertex links are flag simplicial complexes. Moreover, any $CAT(0)$ cubical complex is contractible.
The results below imply that $\widetilde{f}(M(T(A)),w)=P(w+1)$ for some polynomial $P(z)$ with nonnegative coefficients if $T(A)$ is flag.
Lemma 2.35.
(Rowlands, Definition 3.1 on p. 10 and Lemma 3.3 on p. 11 of [29])
Every flag simplicial complex is the crossing complex $\Delta_{P}$ (associated to a $PIP$ $P$) of some $CAT(0)$ cubical complex.
Theorem 2.36.
(Ardila–Owen–Sullivant, Theorem 2.5 on p. 146 of [2], Theorem 2.9 of [29])
The map $P\mapsto\Box_{P}$ is a bijection from the set of $PIP$s to the set of rooted $CAT(0)$ cubical complexes, with inverse given by $\Box\mapsto P_{\Box}$.
Theorem 2.37.
(Rowlands, Theorem 3.5 on p. 13 of [29])
Let $P$ be a $PIP$. Then, $\widetilde{f}(\Box_{P},t)=\widetilde{f}(\Delta_{P},1+t)$.
Input simplicial complexes/cell complexes $A$ such that $T(A)$ is flag yield outputs connected to gamma vectors which arise from $f$-vectors of flag simplicial complexes. It is conjectured that the $h$-vector of every vertex decomposable flag comple is the face vector of some flag complex (Conjecture 3.11 on p. 94 of [11]).